RawlsK29:

how do I solve 2x - 3x/2 = -x + 1/3

RawlsK29:

@Shadow @dude

jhonyy9:

do you think 2x - (3x/2) = -x +(1/3)

RawlsK29:

What do you mean do i think...?

jhonyy9:

in this way wrote correct ?

dude:

\(\large 2x-(\frac{3x}{2})=-x+(\frac{1}{3})\) I think is what jhonny meant